OpenStudy (anonymous):

If θ = 330° and α = 300°, find the exact value of tan θ − sin α.

OpenStudy (mertsj):

\[\tan 330^{o}=\frac{-\sqrt{3}}{3}\] \[\sin 300^{o}=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

so what is \[(-\sqrt{3}/3)-(-\sqrt{3}/2)\]

OpenStudy (mertsj):

\[- \frac{2\sqrt{3}}{6}+\frac{3\sqrt{3}}{6}=\frac{\sqrt{3}}{6}\]